You are testing the null hypothesis that there is no linear relationship between two​ variables, X and Y. From your sample of n = 10 , you determine that r=0.55 .

a. What is the value of t Subscript STATt​?

b. At the alpha=0.05 level of​ significance, what are the critical​ values?
Answer #1

Here' the answer to the question with full concept. Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

n = 10, r = .55

df = n-2 = 10-2 = 8

a. The test statitic for correlation is given by : t = r*sqrt(n-2 / 1-r^2) = .55*sqrt(10-2 / 1-.55^2) = 1.863

b. At alpha = .05, the critical value of t can be had from the t tables. Take df = 8, and alpha = .05, we have t = +/- 2.31

Since our test statistic is between +/- 2.31 we conclude that our correlation is not significant